OpenStudy (anonymous):

Simplify. 4 + 3(10 - 23) A) 8 B) 10 C) 11 D) 12

OpenStudy (anonymous):

None of those answers are correct if you use the order of operations/ PEMDAS

OpenStudy (anonymous):

Please make sure you didn't make a type o :)

OpenStudy (anonymous):

I didn't this is my problem

OpenStudy (anonymous):

It dosn't make scence :(

OpenStudy (anonymous):

Oh. No. Using the order of operations you do what ever is in the parenthesis first then multiply that by 3 THEN add 4. You don't add three and four first.

OpenStudy (anonymous):

I know that

OpenStudy (anonymous):

The answer would be -55

OpenStudy (anonymous):

I use PEMDAS but it doesn't make scence

OpenStudy (dr0zier99):

@Bamawam you meen -35

OpenStudy (anonymous):

Hold on... Let me check my work.

OpenStudy (anonymous):

oh wait also the three is an exponet

OpenStudy (anonymous):

\[4 + 3(10 - 2^{3}\]

OpenStudy (anonymous):

4+3(10-23) 4+3(-13) 4+(-39) -35 I made a type o when I put -55 LOL

OpenStudy (dr0zier99):

do you mean the question is 4^3(10-23)

OpenStudy (dr0zier99):

your answer is 10

OpenStudy (anonymous):

Then the answer is 10.

OpenStudy (anonymous):

how is it 10?

OpenStudy (anonymous):

2^3 is basically 2*2*2 which equals 8. 10 - 8=2 3*2=6 6+4= 10

OpenStudy (anonymous):

oh ok thank you